Find the X value for the vertex : (show all work)
1. f(x)=x^2+4x+6
2.f(x)=2(x-4)^2+5
3.f(x)=-2(x+4)^2+3
4.f(x)=-12x^2+12x+3
1
Expert's answer
2019-07-25T14:02:52-0400
"f(x)=x^2+4x+6"

Solution:


"f(x)=(x+2)^2+2"



"(x+2)^2=0; x+2=0;x=-2;""x=-2 => 0+2=y;\space y=2"

Answer: "(-2,2);"

2.


"f(x)=2(x-4)^2+5"

Solution:

"2(x-4)^2=0;\space x-4=0;\space x=4""x=4 => 0+5=y;\space y=5"

Answer: "(4,5);"

3.

"f(x)=-2(x+4)^2+3"

Solution:

"-2(x+4)^2=0;\space x+4=0;\space x=-4"


"x=-4 => 0+3=y;\space y=3"

Answer:"(-4,3);"

4.

"f(x)=-12x^2+12x+3"

Solution:

"f(x)=-3(2x-1)^2+6;"

"-3(2x-1)^2=0; \space2x-1=0;\space x=\frac{1}{2};"

"x=\frac{1}{2}=>\space 0+6=y;\space y=6"

Answer:"(\frac{1}{2},6);"